The life of a certain AAA batteries is normally distributed with a variance of 100 hours and mean of 550 hours. Find the probability that a battery chosen at random will last no more than 568 hours
Answer #1

Given,

= 550 , = sqrt(100) = 10

We convert this to standard normal as

P( X < x) = P( Z < x - / )

P( X < 568) = P( Z < 568 - 550 / 10)

= P( Z < 1.8)

= 0.9641